# Diverging network architecture of the C. elegans connectome and signaling network

### v1

@@ Abstract
The connectome describes the complete set of synaptic contacts through which neurons communicate. While the architecture of the C. elegans connectome has been extensively characterized, much less is known about the organization of causal signaling networks arising from functional interactions between neurons. Understanding how effective communication pathways relate to or diverge from the underlying structure is a central question in neuroscience. Here, we analyze the modular architecture of the C. elegans signal propagation network, measured via calcium imaging and optogenetics, and compare it to the underlying anatomical wiring measured by electron microscopy. Compared to the connectome, we find that signaling modules are not aligned with the modular boundaries of the anatomical network, highlighting an instance where function deviates from structure. An exception to this is the pharynx which is delineated into a separate community in both anatomy and signaling. We analyze the cellular compositions of the signaling architecture and find that its modules are enriched for specific cell types and functions, suggesting that the network modules are neurobiologically relevant. Lastly, we identify a "rich club" of hub neurons in the signaling network. The membership of the signaling rich club differs from the rich club detected in the anatomical network, challenging the view that structural hubs occupy positions of influence in functional (signaling) networks. Our results provide new insight into the interplay between brain structure, in the form of a complete synaptic-level connectome, and brain function, in the form of a system-wide causal signal propagation atlas.
@@ Introduction
How the structure of a biological system supports its functional repertoire is a central question in biology. A defining feature of this question is that it seeks to relate two lenses into the same biological system. In neuroscience, the relevant structural lens is the network map of synaptic connections in the brain, called a “connectome”, which is typically measured using electron microscopy and captures the wiring of the nervous system. The functional lens is given by network maps that report how the activity of one neuron in the brain either relates to or influences the activity of others. These can be measured by functional MRI, electrophysiology, or calcium imaging. Network science provides a powerful toolkit to characterize and distill key properties of these networks [1, 2]. There has been an implicit hypothesis that since network properties capture fundamental features of the brain they should persist across both of these lenses. However, in studies in humans, there is evidence that network properties can differ, sometimes dramatically, between function and structure, including measures of centrality [3, 4] and community structure [5, 6, 7]. Therefore, the extent to which properties of a brain’s structural network match properties of its functional network remains unclear, especially at brain scale and cellular level.
With few exceptions [8, 9, 10], the link between the synapse level wiring of a brain’s connectome and its function has not been explored in detail and at scale, and the mapping between the two may not be trivial. One reason that structure and function may differ is that some wired connections may exist without a corresponding functional connection. Gene expression or neuromodulators may transiently activate subsets of connections from a menu of all possibilities [11, 12]. Similarly, there is evidence of functional connections that exist in the absence of wiring. Some of these can be attributed to extrasynaptic signaling, where one neuron releases peptides or other transmitters via dense core vesicles that diffuse through the extracellular milieu to activate other neurons [13, 14, 15, 16, 17, 18, 19, 20]. Other functional connections that lack direct wiring can arise from multi-step pathways–so-called “poly-synaptic paths”, where one neuron signals to another, which in turn signals to another [21, 22]. All of these effects make it challenging to translate between the networks of structure and function.
Until now, it has been difficult to empirically compare brain structure and function, as this requires both synapse-level maps of anatomical wiring as well as corresponding causal maps of how one neuron’s activity influences another’s. Recently, we measured a brain-wide causal “Signal Propagation Atlas” of the nematode C. elegans via direct optogenetic activation and simultaneous whole-brain calcium imaging [14]. This directed network describes the flow of signals between 23,433 pairs of neurons in response to controlled and systematic stimulation of individual neurons. Its connections, therefore, represent the “effective” connection between pairs of neurons, taking into account mono- and poly-synaptic signaling pathways, and extrasynaptic signaling. Uniquely and in contrast to previous correlation-based measures of functional connectivity, this dataset provides causal information about signaling between neurons. Critically, this signaling network is accompanied by a complete connectome of chemical synapses and gap junctions between pairs of neurons [23, 24, 25] allowing us to directly compare the physical wiring of a nervous system with the causal signal flow at scales ranging from an individual synapse to the entire brain.
Here, we investigate and compare the properties of both networks: wiring and signaling, focusing on their community structure and rich club. We show that along these dimensions, and with only a few notable exceptions, properties of the signaling network differ from the underlying connectome. We focus further on the hierarchical community structure of the signaling network and show that, despite exhibiting only a weak correspondence with connectome communities, its communities are significantly enriched for specific cell classes and functions. One area of agreement is the community of pharyngeal neurons, which is defined similarly in both networks, but which interacts with the rest of the network differently. Finally, we compare the rich clubs of both networks and find only minimal overlap in terms of membership. Collectively, our findings suggest that the effective signaling network, which measures the collective influence of direct, indirect, and extrasynaptic signaling pathways, has different network properties from the underlying connectome.
@@ Results
Here, we compare the network properties of the C. elegans connectome [23, 24, 25] –the network of chemical and electrical synapses of the brain– to a recently-constructed network map of neural signaling [14] of the genetically identical brain. While the connectome encodes only direct synaptic connections between neurons, the signaling network captures effective connections that arise from the combined effect of not only direct connections but also all interactions that span multiple hops through the anatomical network (poly-synaptic chains), as well as extrasynaptic signaling that is not constrained to flow along synapses. We consider binarized directed connectome and signaling networks of the same head neurons ($N=188$ head neurons from a total of 302 neurons in the entire nervous system) but each having different numbers of edges (corresponding to $M=1151$ for signaling and $M=3250$ for the connectome). Thus, the networks differ in terms of their density ($d_{s}=0.033$ and $d_{a}=0.092$) and their distribution of motifs (Supplementary Fig. 2). The signaling network has an average clustering coefficient of $0.13$, a bit less than half of the anatomical network’s clustering coefficient of $0.3$. The C. elegans connectome is well known to be a small world network [26, 27]. Similarly, we find that the signaling network is also small world ($\sigma_{\text{s}}=4.76$ and $\sigma_{\text{a}}=2.52$ for the signaling and anatomical head networks, respectively).
@@
We chose to compare the complex architecture of wiring and signaling networks by means of their hierarchical modular structure. Neural networks are hierarchically divided into sub-networks, referred to as modules, communities, or clusters [28, 29, 30, 31, 32, 33]. This modular composition is thought to shape neural communication, by providing a balance between segregation into specialized functional units and integration of information across distant neural elements [7] and therefore informs how the nervous system performs computations[34, 35, 36]. In fMRI studies, abnormalities in modular structure serve as biomarkers of neuropsychiatric conditions [37]. Modular structure has been used to interpret both anatomical [38, 39, 7] and functional [5, 40] representations of neural networks. Yet, it remains unclear whether modular structure should be conserved in these different views, and in some cases evidence suggests they are not [6, 41].
We explicitly tested whether the anatomical and signaling descriptions of the C. elegans brain shared a similar modular structure. We used a hierarchical variant of the stochastic block model (SBM) to independently subdivide the wiring and signaling networks into nested communities [42, 43, 44, 45, 46]. We found stark differences in the modular organization of these two views into the C. elegans brain. The anatomical connectome was clustered into 20 communities in the first hierarchical level, with the smallest containing only one neuron and the largest containing 20 (Fig. 1 b, anatomical clusters are denoted by Arabic numerals). In contrast, the signal propagation network was optimally partitioned into six communities in the first hierarchical level, the smallest of which contained 17 neurons and the largest of which contained 47 neurons (Fig. 1 a, signal propagation clusters are denoted by Roman numerals). Interestingly, the signaling network exhibited fewer hierarchical levels compared to the anatomical network (three levels composed of six, three, and one communities for the signaling network, and five levels composed of twenty, eight, three, two, and one communities for the connectome).
@@
Next, we investigated the distribution of bilateral pairs across communities. A majority of C. elegans neurons consist of bilaterally symmetric pairs that are often connected to each other via gap junctions and tend to have similar gene expression [15], wiring [23], and correlated activity patterns [9]. As we would expect, bilaterally symmetric neurons are much more likely to be found in the same communities than across different communities, for both the anatomical and signaling networks. In the anatomical network, bilateral neuron pairs were 38 times more likely to be in the same community than in different communities ($p=1.5\times 10^{-210}$) and in the signaling network, they were twice as likely ($p=4\times 10^{-45}$). This finding is consistent with the hypothesis that some features should be conserved across both network descriptions.
@@
If the community architecture of the signaling and anatomical networks is similar we would expect signal propagation communities to either correspond to specific partners in the anatomical network, or since there are fewer signaling communities, we would expect some signaling communities to entirely contain anatomical communities. To test whether signaling and anatomical communities follow this straightforward correspondence, we first calculated the pairwise overlap between each of the signal propagation and anatomical communities (Fig. 2 a). With only one exception (community I), signal propagation communities did not neatly map onto or wholly contain anatomical communities. Most signal propagation communities had neuron membership that was spread across multiple anatomical communities (rows of Fig. 2 a). Similarly, all anatomical communities had neuron membership that was spread across many signal propagation communities (columns of Fig. 2 a). These findings suggest that most signal propagation communities are not trivially mapped onto nor trivially contain anatomical communities.
Even without a trivial mapping between signal propagation and anatomical communities, we nonetheless sought to identify the “best-matched” signal-anatomy community pairs. We therefore calculated the extent to which each signal propagation community was enriched for anatomical communities, and compared this to a null hypothesis in which neurons are assigned to communities at random [47], (Fig. 2 b). Clustering was performed on this enrichment map to assign best-matched pairs (Fig. 2 c), as described in the methods. The majority of signaling communities (four of six) were statistically significantly enriched for at least one anatomical community. However, to be a relatively strong match, we would expect a signaling community to have statistically significant enrichment for only one community. We find only two of the signaling communities had statistically significant enrichment for exactly one of the anatomical communities. This suggests that the majority of signaling communities (four of six) had only weak pairing to their best-matched anatomical pair.
Across all analyses, the pairing of signaling community I and anatomical community 20 stands out as an exception. These two communities neatly correspond to one another– 93% of signaling community I’s neurons are members of anatomical community 20, while 75% of anatomical community 20’s neurons are members of signaling community I. Similarly, both communities are statistically significantly enriched only for one another.
@@
Why might communities I and 20 neatly correspond, when others do not? These two communities are comprised almost exclusively of pharyngeal neurons (Fig. 3 c,d). The pharynx has long been hypothesized to be largely functionally autonomous [48, 49], partly because it makes only two extrapharyngeal synaptic connections to the rest of the network (Fig. 1 a) [23, 24, 25]. Its role in controlling pharyngeal pumping and feeding behaviors has historically been assumed to operate largely independently of the rest of the network [49, 48]. This stark anatomical isolation may explain why both the structural and signaling networks preserve this community organization. It is interesting that, even with this direct correspondence between the anatomical and signaling communities of the pharyngeal neurons, in the signaling network, the community of pharyngeal neurons is much more highly connected to the other communities than it is in the anatomical network (Fig. 1 a), highlighting differences between the brain’s wiring and signaling.
@@
Collectively, our results suggest that, in general, modular structure is not conserved across anatomical and functional descriptions of the brain. In the discussion, we speculate that a combination of polysynaptic signaling, extrasynaptic signaling, and non-functional anatomical wiring may contribute to these differences. The pharynx is an exception and suggests that in cases where a sub-network of neurons is extremely isolated in the anatomical network, its modularity may be preserved even in the functional description. However, even here, it is notable that the patterns of connectivity between the modules still differ.
@@
Clustering of anatomical networks has previously been used to identify communities of neurons that are enriched for cell types or known functional cellular roles, including in Drosophila and C. elegans [50, 51, 52, 53, 54, 55]. Consistent with prior findings, our modular clustering of the C. elegans anatomical network recapitulates anatomical communities that are each significantly enriched for no more than one of the three C. elegans neuron cell types classes: Sensory, Inter- and Motor neurons [27](Fig. 3 c; Supplementary Fig. 4 a; Table 2). Similarly, they are each enriched for no more than one of five predefined cell roles (pharyngeal, amphid sensory, locomotory inter-, vnc motor and ring interneurons) (Fig. 3 d; Supplementary Fig. 4 b; Table 2). Enrichment is calculated by comparison to a null hypothesis of randomly shuffled assignments as described in methods. We next explored cell-type annotation in the signaling network.
Signaling network communities, like the anatomical communities, also showed clear enrichment for distinct cell types. Each signaling community was statistically significantly enriched for no more than one of three previously defined cell type classes (Fig. 3 a,b; Supplementary Fig. 4 a; Table 1). Only one of the six communities, V, showed no significant enrichment for any cell type.
Similarly, cell roles are well segregated into different signaling communities. As mentioned in the previous section community I contains almost all the pharyngeal neurons. All but one locomotory interneuron are members of community VI (Fig. 3 c; Supplementary Fig. 4 b; Table 1). Along with interneurons involved in locomotion, community VI was also enriched for the amphid sensory neurons. The amphid sensory neurons are the primary olfactosensory neurons of C. elegans and play a role in chemotaxis and mechanosensation [56, 57, 58]. The locomotor interneurons AVA, AVD, and AVE from this community, among others, translate this sensory signal into behavior to control and initiate backward locomotion [59, 60, 61, 62, 63]. Similarly, neuron AVB from this community is a key part of the forward locomotory circuit [64]. The community structure may therefore have defined a cluster in such a way as to capture a sensorimotor pathway.
@@
It is commonly expected that there should be overall feed-forward signaling from sensory neurons to interneurons to motor neurons [27]. Anatomically, synaptic edges are enriched in a feed-forward manner from sensory neurons to interneurons to motor neurons (Supplementary Fig. 5 b,c). By contrast, the signaling network is not appreciably feed-forward along this path compared to a shuffle control (Supplementary Fig. 5 b,c). This is also visible in the connectivity between and among communities. We do not see any obvious evidence of feed-forwardness when looking at how communities that are enriched for sensory, inter, and motor neurons signal to each other (Supplementary Fig. 6).
@@
To gain insight into large-scale network organization, we identified stereotyped inter-community relations, called “community motifs” that assign a label to how each community interacts with each other community, either assortative, disassortative, or core-periphery (Fig. 4 a) [67]. Different community interactions provide hypotheses about the roles that communities may play when interacting. For example, the presence of assortative communities, which are internally dense but externally sparse, can be evidence of separate computational units. We observe, as expected, that the majority of community interactions in both signal propagation and anatomical networks are assortative ($70\%$ for both networks) [67, 50, 68]. In both networks, neurons most frequently participated in assortative interactions Fig. 4 b).
We were particularly interested in investigating core-periphery interactions because they can be evidence of broadcasting and integrating information. Core-periphery interactions are comprised of one densely connected core of “driver” nodes with substantial connections to a periphery that, in turn, is sparsely connected internally. We observed that in the anatomical network, $26\%$ and in the signaling network, $30\%$ of community interactions are core-periphery and we wondered which types of neurons may be involved in this broadcasting and integrating motif.
@@
As participation in the core of a core-periphery interaction could potentially indicate a role for broadcasting or integrating signals, we were curious if the pharyngeal and amphid neurons fit one of these two roles. By inspecting the number of ingoing and outgoing edges in the signaling network, we find that pharyngeal neurons show more broadcasting-like behavior perhaps suggesting a role of sending pumping or food-related signals broadly to the rest of the nervous system (Fig. 4 c). Counterintuitively amphid sensory neurons have more incoming than outgoing edges in the signaling network, possibly suggesting that more research is needed to explore whether they could be serving as integrators (Fig. 4 c). The ratios of incoming to outgoing edges could not trivially be explained by the number of times in which amphid or pharyngeal neurons were observed or stimulated (Supplemental Fig. 8 a,b,c,d).
This change in participation in interaction styles of the same group of neurons between the two networks was also observed in other cell roles, including locomotory interneurons which are enriched for participation in disassortative interactions in anatomy, a motif that signifies a capacity for transmitting signals across the boundaries of communities and which is entirely absent from the signaling network. This switch in interaction styles for the different neuron types between the anatomical network and the signaling network suggests that signaling may be performing neural processing in ways different from what we may expect from looking solely at the anatomical network topology.
@@
Like other complex networks [69, 70], the C. elegans connectome exhibits “rich-club” connectivity [71], in which highly connected hub nodes are also densely connected to each other. We wondered whether we should also expect to see rich-club features in the signaling network.

### v2

@@ Abstract
The connectome describes the complete set of synaptic contacts through which neurons communicate. While the architecture of the C. elegans connectome has been extensively characterized, much less is known about the organization of causal signaling networks arising from functional interactions between neurons. Understanding how effective communication pathways relate to or diverge from the underlying structure is a central question in neuroscience. Here, we analyze the modular architecture of the C. elegans signal propagation network, measured via calcium imaging and optogenetics, and compare it to the underlying anatomical wiring measured by electron microscopy. Compared to the connectome, we find that signaling modules are not aligned with the modular boundaries of the anatomical network, highlighting an instance where function deviates from structure. However, we find that some of the most striking features of the anatomical network are preserved, as exemplified by the pharynx, which is delineated into a separate community in both anatomy and signaling. We analyze the cellular compositions of the signaling architecture and find that its modules are enriched for specific cell types and functions, suggesting that the network modules are neurobiologically relevant. Lastly, we identify a "rich club" of hub neurons in the signaling network. The membership of the signaling rich club differs from the rich club detected in the anatomical network, challenging the view that structural hubs occupy positions of influence in functional (signaling) networks. The only overlap between the two rich clubs is given by neurons AVEL/R, which have some of the highest degrees in the anatomical network, again illustrating the preservation of the most pronounced features of the network. Our results provide new insight into the interplay between brain structure, in the form of a complete synaptic-level connectome, and brain function, in the form of a system-wide causal signal propagation atlas.
@@ Introduction
How the structure of a biological system supports its functional repertoire is a central question in biology. A defining feature of this question is that it seeks to relate two lenses into the same biological system. In neuroscience, the relevant structural lens is the network map of synaptic connections in the brain, called a “connectome”, which is typically measured using electron microscopy and captures the wiring of the nervous system. The functional lens reports how the activity of one neuron in the brain either relates to or influences the activity of others. These can be measured by functional MRI, electrophysiology, or calcium imaging. In neuroscience, going back to Cajal [1], there is a strong expectation that structure should inform function [2, 3, 4].
Network science provides a powerful toolkit to characterize and distill key properties of networks [5, 6] that, in some cases, can help link structure to function. For example, in a simple model of disease spread, the characteristic path-length of the static network of human interactions (structure) has a simple scaling relation with the time to global infection (function) [7, 8].
The success of examples like these has led to a common implicit hypothesis: that since network properties capture fundamental features of a network, they might persist across network descriptions of both the structural and functional lenses of the brain. This hypothesis is explicit in neuroscience investigations into how anatomy relates to fMRI or EEG recordings [9, 10]. It also motivates network investigations into connectomes of the simple nematode C. elegans [11, 12, 13, 14, 15] and of Drosophila [16, 17].
However, it remains unclear whether features of the network architecture should be preserved in both a network description of a brain’s connectome and a network description of its neural function or dynamics. Many common network features came about because they are useful in the context of specific dynamical assumptions, e.g., small worldness is particularly informative for describing systems with simple linear dynamics, such as disease transmission or random walks [8]. However, the brain has more complex nonlinear dynamics, and there, the applicability of many of these metrics is less obvious [8]. Similarly, other properties of dynamics can lead to differences in network descriptions of structure and function. For example, if neural signals rapidly propagate deep through the network across multi-hop paths, such “polysynaptic” signaling would contribute towards differences in features of the structural and functional networks [10, 18]. While studies in humans are inherently limited in their scale or resolution, there is evidence that network properties can differ, sometimes dramatically, between network descriptions of correlations in BOLD activity and in gross anatomy, including measures of centrality [19, 20] and community structure [21, 22, 23].
The extent of agreement between network descriptions of a connectome and neural activity may also depend on how well the wires of the connectome themselves capture all the relevant information for governing neural dynamics. While synaptic transmission is clearly important for neural signaling, there may be other factors not visible from the wires that also contribute. For example, wired connections may exist without a corresponding functional connection because key receptors are absent or downregulated. Gene expression or neuromodulators may alter the number, mix, or physiological properties of receptors or transmitters to transiently activate subsets of wired connections from a menu of all possibilities [24, 25]. Similarly, there is evidence of functional connections that exist in the absence of wiring. Some of these can be attributed to extrasynaptic signaling, where one neuron releases peptides or other transmitters via dense core vesicles that diffuse through the extracellular milieu to activate other neurons and therefore would not be visible in wiring [26, 27, 28, 29, 30, 31, 32, 33].
With few exceptions [34, 15, 14], the link between the synapse-level wiring of a brain’s connectome and its function has not been explored in detail and at scale. Therefore, the extent to which network properties of a brain’s connectome match those of a network description of its dynamics remains unclear, especially at brain scale and cellular level. One challenge has been a lack of synapse-level maps of anatomical wiring as well as corresponding causal maps of how one neuron’s activity influences another’s.
Recently, we measured a brain-wide causal “Signal Propagation Atlas” of the nematode C. elegans via direct optogenetic activation and simultaneous whole-brain calcium imaging [27]. Critically, this signaling network is accompanied by a complete connectome of chemical synapses and gap junctions between pairs of neurons [11, 35, 36]. This gives us the unique opportunity to directly compare network features of the physical wiring of a nervous system with its causal signal flow empirically for the first time at this scale and resolution.
Here, we investigate and compare the properties of both networks: wiring and signaling, focusing on their community structure and rich club. We show that for these network properties, only the most striking features are preserved, and otherwise, there are significant differences between the two networks. For example, both networks have a distinct community corresponding to the pharynx, but most other communities do not correspond across the two networks. Similarly, the most connected member of the rich club of the signaling network is also the most connected member of the connectome, but there is little other overlap in rich club membership.
Collectively, our findings suggest that the effective signaling network, which measures the collective influence of direct, indirect, and extrasynaptic signaling pathways, has different network properties from the underlying connectome.
@@ Results
Here, we compare the network properties of the C. elegans connectome [11, 35, 36] –the network of chemical and electrical synapses of the brain– to a recently-constructed network map of neural signaling [27] of the genetically identical brain. While the connectome encodes only direct synaptic connections between neurons, the signaling network captures effective connections that arise from the combined effect of not only direct connections but also all interactions that span multiple hops through the anatomical network (poly-synaptic chains), as well as extrasynaptic signaling that is not constrained to flow along synapses. We consider binarized directed connectome and signaling networks of the same head neurons ($N=188$ head neurons from a total of 302 neurons in the entire nervous system), but each having different numbers of edges (corresponding to $M=1151$ for signaling and $M=3250$ for the connectome). Binarization of the anatomical network was performed by counting the average number of connections across the four connectomes [11, 36] (see Methods). By default, a single synaptic contact in any of the four connectomes was sufficient for inclusion as an edge in the network. For the signaling network, pairs of neurons that exhibited a calcium response with a q-value of less than 0.05 were included. Thus, the networks differ in terms of their density ($d_{s}=0.033$ and $d_{a}=0.092$) and their distribution of motifs (Extended Data Fig. 1). The signaling network has an average clustering coefficient of $0.13$, a bit less than half of the anatomical network’s clustering coefficient of $0.3$. The C. elegans connectome is well known to be a small world network [7, 12]. Similarly, we find that the signaling network is also small world ($\sigma_{\text{s}}=4.76$ and $\sigma_{\text{a}}=2.52$ for the signaling and anatomical head networks, respectively).
@@
We chose to compare the complex architecture of wiring and signaling networks by means of their hierarchical modular structure. Neural networks are hierarchically divided into sub-networks, referred to as modules, communities, or clusters [37, 38, 39, 40, 41, 42]. This modular composition is thought to shape neural communication by providing a balance between segregation into specialized functional units and integration of information across distant neural elements [23] and therefore informs how the nervous system performs computations[43, 44, 45]. In fMRI studies, abnormalities in modular structure serve as biomarkers of neuropsychiatric conditions [46]. Modular structure has been used to interpret both anatomical [47, 48, 23] and functional [21, 49] representations of neural networks. Yet, it remains unclear whether modular structure should be conserved in these different views, and in some cases, evidence suggests they are not [22, 50].
We explicitly tested whether the anatomical and signaling descriptions of the C. elegans brain shared a similar modular structure. We used a hierarchical variant of the stochastic block model (SBM) to independently subdivide the wiring and signaling networks into nested communities [51, 52, 53, 54, 55, 56]. The anatomical and signaling networks differ in the number of communities and in their hierarchical organization. The anatomical network had more communities, it was optimally partitioned into 20 communities in the first hierarchical level, with the smallest containing only one neuron and the largest containing 20 (Fig. 1 b, anatomical clusters are denoted by Arabic numerals). In contrast, the signal propagation network had fewer and was optimally partitioned into six communities in the first hierarchical level, the smallest of which contained 17 neurons and the largest of which contained 47 neurons (Fig. 1 a, signal propagation clusters are denoted by Roman numerals). The larger number of communities in the organization of the anatomical network was not trivially explained by its higher density. We performed two comparisons using different versions of the signaling and anatomy networks, in which we enforced the network density to be the same by adjusting the threshold for inclusion of edges of either network, called “enforced density controls.” In one case, we compared the existing signaling network to a more conservative anatomical network in which we required neurons to have more synaptic contacts for inclusion. In another case, we compared the existing anatomical network to a more permissive signaling network that included lower signal-to-noise responses. In both sets of comparisons, the anatomical network had more communities (Extended Data Fig. 12, 13). This difference was also robust to the use of a more stringent thresholding– a “conservative control” in which both had more restrictive inclusion criteria but different densities (Extended Data Fig. 11).
@@
Next, we investigated the distribution of bilateral pairs across communities. A majority of C. elegans neurons consist of bilaterally symmetric pairs that are often connected to each other via gap junctions and tend to have similar gene expression [28], correlated activity patterns [15], and often (but not always [57]) similar wiring [11]. As we would expect, bilaterally symmetric neurons are much more likely to be found in the same communities than across different communities, for both the anatomical and signaling networks. In the anatomical network, bilateral neuron pairs were 38 times more likely to be in the same community than in different communities ($p=1.5\times 10^{-210}$) and in the signaling network, they were twice as likely ($p=4\times 10^{-45}$). Even when adjusting thresholds for inclusion in the networks, this trend persisted: two bilaterally symmetric neurons were more likely to be clustered in the same versus different communities, and the degree of enrichment was higher for anatomy than for signaling (Extended Data Table 1). This finding is consistent with the hypothesis that some features should be conserved across both network descriptions.
@@
If the community architecture of the signaling and anatomical networks is similar, we would expect signal propagation communities to either correspond to specific partners in the anatomical network, or, since there are fewer signaling communities, we would expect some signaling communities to entirely contain anatomical communities. To test whether signaling and anatomical communities follow this straightforward correspondence, we first calculated the pairwise overlap between each of the signal propagation and anatomical communities (Fig. 2 a). Signal propagation community I had large overlap with Anatomical Community 20 both of which correspond to the pharyngeal subnetwork. Other than the pharynx, signal propagation communities did not neatly map onto or wholly contain anatomical communities (adjusted rand score $ARS=0.105$). Most signal propagation communities had neuron membership that were spread across multiple anatomical communities (rows of Fig. 2 a). Similarly, most anatomical communities had neuron membership that was spread across many signal propagation communities (columns of Fig. 2 a).
We further tested whether the amount of overlap we do observe could be explained by chance. We therefore calculated the extent to which each signal propagation community was enriched for anatomical communities, and compared this to a null hypothesis in which neurons are assigned to communities at random [58], (Fig. 2 b). Of all possible signal-anatomy community pairings only seven were significantly enriched, considering multiple hypothesis testing. Overlap was not improved by altering the stochastic block model algorithm to force it to partition each network into the same number of communities ($ARS_{same6}=0.002$ and $ARS_{same20}=0.014$; Extended Data Fig. 4).
These findings suggest that most signal propagation communities are not trivially mapped onto nor trivially contain anatomical communities. The pairing of signaling community I and anatomical community 20 stands out as an exception. The correspondence between pharyngeal communities in signaling and anatomy persists even in our conservative and enforced density controls (Extended Data Fig. 11, 12, 13).
@@
Why might pharyngeal communities neatly correspond, when others do not? The pharynx has long been hypothesized to be largely functionally autonomous [59, 60], partly because it makes only two extrapharyngeal synaptic connections to the rest of the network (Fig. 1 a) [11, 35, 36]. Its role in controlling pharyngeal pumping and feeding behaviors has historically been assumed to operate largely independently of the rest of the network [60, 59]. This stark anatomical isolation may explain why both the structural and signaling networks preserve this community organization. It is interesting that, even with this direct correspondence between the anatomical and signaling communities of the pharyngeal neurons, in the signaling network, the community of pharyngeal neurons is much more highly connected to the other communities than it is in the anatomical network (Fig. 1 a), highlighting differences between the brain’s wiring and signaling.
@@
Collectively, our results suggest that, while the most striking communities may be preserved, in general, modular structure is not conserved across anatomical and functional descriptions of the brain. In the discussion, we speculate that a combination of polysynaptic signaling, extrasynaptic signaling, and non-functional anatomical wiring may contribute to these differences. The pharynx suggests that in cases where a sub-network of neurons is extremely isolated in the anatomical network, its modularity may be preserved even in the functional description. However, even here, it is notable that the patterns of connectivity between the modules still differ.
@@
Clustering of anatomical networks has previously been used to identify communities of neurons that are enriched for cell types or known functional cellular roles, including in Drosophila and C. elegans [16, 61, 62, 63, 64, 65]. Consistent with prior findings, our modular clustering of the C. elegans anatomical network recapitulates anatomical communities that are each significantly enriched for no more than one of the three C. elegans neuron cell types classes: Sensory, Inter- and Motor neurons [12](Fig. 3 c; Extended Data Fig. 5 a; Table 2). Similarly, they are each enriched for no more than one of five predefined cell roles (pharyngeal, amphid sensory, locomotory inter-, vnc motor and ring interneurons) (Fig. 3 d; Extended Data Fig. 5 b; Table 2). Enrichment is calculated by comparison to a null hypothesis of randomly shuffled assignments as described in methods. We next explored cell-type annotation in the signaling network.
Signaling network communities, like the anatomical communities, also showed clear enrichment for distinct cell types. Each signaling community was statistically significantly enriched for no more than one of three previously defined cell type classes (Fig. 3 a,b; Extended Data Fig. 5 a; Table 1). Only one of the six communities, V, showed no significant enrichment for any cell type.
Similarly, cell roles are well segregated into different signaling communities. As mentioned in the previous section community I contains almost all the pharyngeal neurons. All but one locomotory interneuron are members of community VI (Fig. 3 c; Extended Data Fig. 5 b; Table 1). Along with interneurons involved in locomotion, community VI was also enriched for the amphid sensory neurons. The amphid sensory neurons are the primary olfactosensory neurons of C. elegans and play a role in chemotaxis and mechanosensation [66, 67, 68]. The locomotor interneurons AVA, AVD, and AVE from this community, among others, translate this sensory signal into behavior to control and initiate backward locomotion [69, 70, 71, 72, 73]. Similarly, neuron AVB from this community is a key part of the forward locomotory circuit [74]. The community structure may therefore have defined a cluster in such a way as to capture a sensorimotor pathway.
@@
It is commonly expected that there should be overall feed-forward signaling from sensory neurons to interneurons to motor neurons [12]. Anatomically, synaptic edges are enriched in a feed-forward manner from sensory neurons to interneurons to motor neurons (Extended Data Fig. 6 b,c). By contrast, the signaling network is not appreciably feed-forward along this path compared to a shuffle control (Extended Data Fig. 6 b,c). This is also visible in the connectivity between and among communities. We do not see any obvious evidence of feed-forwardness when looking at how communities that are enriched for sensory, inter, and motor neurons signal to each other (Extended Data Fig. 7).
@@
To gain insight into large-scale network organization, we identified stereotyped inter-community relations, called “community motifs” that assign a label to how each community interacts with each other community (Fig. 4 a) [77]. Unlike some common methods for community detection designed to detect “assortative” community structure (for example, modularity maximization [78] or infomap [79]), the stochastic block model can detect generic types of community structure. Specifically, given two communities $r$ and $s$, we considered all community pairs and their within- ($w_{rr}$ and $w_{ss}$, given by the number of edges between nodes in the community divided by the number of possible edges between nodes in the community) and between-community densities ($w_{rs}$ given by the number of edges between nodes in community $r$ and community $s$ divided by the number of possible edges between the two communities), classifying the pairwise interactions as one of the three different community motifs (Figure 4 a). Two communities are classified as assortative if $w_{rr}>w_{rs}$ and $w_{ss}>w_{rs}$. An interaction between communities is disassortative if $w_{rr}<w_{rs}$ and $w_{ss}<w_{rs}$. Lastly, two communities are classified as core-periphery in the case that $w_{rr}>w_{rs}>w_{ss}$.
Different community interactions provide hypotheses about the roles that communities may play when interacting. For example, the presence of the internally dense but externally sparse assortative communities, has been thought to be evidence of separate computational units. We observe, as expected, that the majority of community interactions in both signal propagation and anatomical networks are assortative ($70\%$ for both networks) [77, 16, 80]. In both networks, neurons most frequently participated in assortative interactions (Fig. 4 b).
We were particularly interested in investigating core-periphery interactions because they can be evidence of broadcasting and integrating information. Core-periphery interactions are comprised of one densely connected core of “driver” nodes with substantial connections to a periphery that, in turn, is sparsely connected internally. Here in particular we might expect mismatch between signaling and anatomy because, in contrast to anatomy, signaling reports result of many hops of signals through the network, which result in broadcasting or integrating that wouldn’t be easily identified from the direct connections of the connectome. We observed that in the anatomical network, $26\%$ and in the signaling network, $30\%$ of community interactions are core-periphery and we wondered which types of neurons may be involved in this broadcasting and integrating motif.
@@
As participation in the core of a core-periphery interaction could potentially indicate a role for broadcasting or integrating signals, we were curious if the pharyngeal and amphid neurons fit one of these two roles. By inspecting the number of ingoing and outgoing edges in the signaling network, we find that pharyngeal neurons show more broadcasting-like behavior perhaps suggesting a role of sending pumping or food-related signals broadly to the rest of the nervous system (Fig. 4 c). Counterintuitively amphid sensory neurons have more incoming than outgoing edges in the signaling network, possibly suggesting that more research is needed to explore whether they could be serving as integrators (Fig. 4 c). The ratios of incoming to outgoing edges could not trivially be explained by the number of times in which amphid or pharyngeal neurons were observed or stimulated (Extended Data Fig. 9 a,b,c,d).
This change in participation in interaction styles of the same group of neurons between the two networks was also observed in other cell roles, including locomotory interneurons which are enriched for participation in disassortative interactions in anatomy, a motif that signifies a capacity for transmitting signals across the boundaries of communities and which is entirely absent from the signaling network. This switch in interaction styles for the different neuron types between the anatomical network and the signaling network suggests that the functional role of neurons (broadcaster, integrator, etc) may be hard to predict from network analysis of underlying anatomy.
@@
Like other complex networks [81, 82], the C. elegans connectome exhibits “rich-club” connectivity [83], in which highly connected hub nodes are also densely connected to each other. Hub neurons have been shown to be very relevant for neural dynamics [8] and therefore are a feature where we might expect an overlap between structure and function. We wondered whether we should also expect to see rich-club features in the signaling network, and which neurons would be involved.